what weight of am alloy containing 10% silver must be melted with an alloy containing60% silver to obtain 10lbs of an alloy contain 40% silver
Answer by mananth(16946)

percent ---------------- quantity
Silver Alloy I 10.00% ---------------- x pounds
Silver Alloy II 60.00% ------ 10 - x pounds
Mixture 40.00% ---------------- 10
Total 10 pounds
10.00% x + 60.00% ( 10 - x ) = 40.00% * 10
10 x + 60 ( 10 - x ) = 400
10 x + 600 - 60 x = 400
10 x - 60 x = 400 - -600
-50 x = -200
/ -50
x = 4 pounds 10.00% Silver Alloy I
6 pounds 60.00% Silver Alloy II

m.ananth@hotmail.ca